Wardell Company purchased a mini computer on January 1, 2016, at a cost of $42,550. The computer has been depreciated using the straight-line method over an estimated five-year useful life with an estimated residual value of $4,300. On January 1, 2018, the estimate of useful life was changed to a total of 10 years, and the estimate of residual value was changed to $930.

Required:
1. Prepare the appropriate adjusting entry for depreciation in 2018 to reflect the revised estimate.
2. Prepare the appropriate adjusting entry for depreciation in 2018 to reflect the revised estimate. Assuming that the company uses the sum-of-the-years'-digits method instead of the straight-line method.
Answer #1

1) Original Depreciation = (42550-4300/5) = 7650

Accumulated depreciation = 7650*2 = 15300

Revised depreciation = (42550-15300-930)/8 = 3290

Date account and explanation Debit Credit
Dec 31,2018 Depreciation expense 3290
Accumulated depreciation-Mini Computer 3290
(To record dep)

2)

Accumulated depreciation = (42550-4300)*9/15 = 22950

Revised depreciation = (42550-22950-930)*8/36 = 4149

Date account and explanation Debit Credit
Dec 31,2018 Depreciation expense 4149
Accumulated depreciation-Mini Computer 4149
(To record dep)